# GRINNs: Godunov-Riemann Informed Neural Networks for Learning Hyperbolic Conservation Laws

### v1

@@ 5 Conclusions
GRINNs offer interpretable shallow feedforward NNs for tackling the above problems. Unlike other hybrid SciML methods that approximate the numerical fluxes, GRINNs learn the solution of the Godunov scheme by directly approximating the physical flux functions per se. This approach enhances interpretability, thus respecting explicitly the underlying conservation laws as approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. Moreover, GRINNs provide a versatile framework for identifying flux functionals directly from data. This approach accommodates systems of conservation laws and allows for the inclusion of source terms—whether known or unknown—in the associated hyperbolic PDEs. As we show, GRINNs are able to approximate accurately the closures of four benchmark problems, which exhibit traveling shocks that emerge at finite times. Based on the solution of the inverse problem, which actually approximates the physical flux function required for the solution operator of the Godunov-type scheme, we solved the forward problem for long time periods, showing that GRINNs result to a high numerical approximation accuracy, preserving numerical stability and the conservation laws.
The training and performance of GRINNs, like other SciML methods, benefit from partial knowledge of the physics underlying the missing terms (see section 4.4, where GRINNs converge with the “correct” predictor selection.) If no physical insight is available, then techniques such as symbolic regression [75, 6], manifold learning (e.g., parsimonious diffusion maps) [15, 22, 20, 25, 56], and SINDy [8, 10] can be exploited to learn a parsimonious set of variables that can parametrize the unknown closures. Furthermore, GRINNs can be extended to analyze real-world data derived from microscopic simulations/observations, facilitating the discovery of macroscopic hyperbolic PDEs for the emergent dynamics. This application will further require parameter inference, a task that can be addressed by including physical parameters as additional GRINNs predictors. Finally, using GRINNs to discover macroscopic closures connects directly to control theory, which can be exploited for enabling distributed control in large-scale microscopic models [47, 48], given GRINNs-derived information at the macroscopic level.
@@ Appendix A Benchmark problems and data acquisition
Here, we present the four benchmark problems upon which the performance of GRINNs is evaluated on, namely the inviscid Burgers’ equation, the LWR equation of traffic flow, the SW equations of fluid dynamics, and the PW equations of traffic flow. We additionally provide the details for the solution of the forward problem with the high-resolution Godunov-type FV scheme in section 3.1, which are required for the collection of training/validation/testing data sets.
@@ A.1 Burger’s equation
In spite of the smooth initial conditions, contact discontinuities are formed within the selected time interval, in particular shock waves and rarefactions; the latter are not transonic due to the positive initial conditions. Thus, for the numerical solution of the forward problem via the Godunov-scheme in section 3.1, it is sufficient to employ the Roe solver, which leads to the scalar, in this case, Roe matrix $\hat{a}_{i-1/2}=\bar{u}$, where $\bar{u}=(u_{l}+u_{r})/2$; the derivation is provided in section C.1. As already discussed in section 3.3, the Van-Leer flux-limiter function is employed. The resulting numerical solution of the Burgers’ equation for the Gaussian initial data with $\mu=2$ and $\sigma=0.2$ is depicted in Figure A.1a,b, where it is clearly shown that the initial bump quickly forms a right going shock wave, followed by a rarefaction spread out towards the left; for comparison to the analytic solution, see Chapter 4 of [35]. From the 4 numerical solutions of the Burgers’ equation, we randomly split the 15-15-70% of the time-series data to form the training/validation/testing data sets of the GRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=2$ and $\sigma=0.2$ are denoted with black circles in Figure A.1a.
@@
Figure A.1: Numerical solution of the Burgers’ and LWR equations obtained with the high-resolution Godunov-scheme in section 3.1 for (a,b) $x\in[-1,1]$, $t\in[0,3]$, with the initial data $u(0,x)=2~{}exp(-x^{2}/(2\cdot 0.2^{2}))$ shown in panel (b), and (c,d) $x\in[-20,20]$, $t\in[0,60]$ with the $\rho(0,x)=exp(-(x+10)^{2}/(2\cdot 2^{2}))$ shown in panel (d). Periodic boundary conditions are used for the Burgers’ equation and outflow ones for the LWR equation. The black circles on (a,c) denote the time steps $n_{t}$, collected from the depicted numerical solutions, to form the training set of GRINNs.
@@ A.2 LWR equation of traffic flow
Similarly to the Burgers’ equation, in spite of the smooth initial conditions, contact discontinuities are formed within the selected time interval. For the numerical solution via the Godunov-scheme in section 3.1, it is sufficient to employ the Roe solver. The scalar, for the LWR equation, Roe matrix takes the form $\hat{a}_{i-1/2}=v_{max}(1-2\bar{\rho})$, where $\bar{\rho}=(\rho_{l}+\rho_{r})/2$; see the derivation in section C.2. The resulting numerical solution of the LWR equation, using the Van-Leer flux-limiter function, for the Gaussian initial conditions with $\sigma=2$ is depicted in Figure A.1c,d. In contrast to the Burgers’ equation, the non-linear flux term of the LWR equation $f(\rho)=v_{max}\rho(1-\rho)$ is not convex, but concave. Because of that, as shown in Figure A.1c, the initial smooth profiles quickly form a right going shock wave, which is now following the right-spreading rarefaction (instead of preceding it, as shown in Figure A.1a for the Burgers’ equation). This behavior is well aligned with the relaxation of traffic jams reported for agent-based car-following models; see Chapter 11 of [41] for a direct comparison. From the 4 numerical solutions of the LWR equation, we randomly split the 15-15-70% of the time-series data to form the training/validation/testing data sets of the GRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=1$ and $\sigma=2$ are denoted with black circles in Figure A.1c.
@@ A.3 The SW equations
the latter being the so-called Roe average. For the derivation of $\hat{\mathbf{A}}_{i-1/2}$, which is presented in detail in section C.3, we followed the general approach introduced by [66], resulting in the same Roe matrix as the ones derived in [41, 35]. The resulting numerical solution of the SW equations, using the Van-Leer flux-limiter function, for the Gaussian initial data with $\sigma=0.4$ is depicted in Figure A.2a,b,c for $h$ and $q$, respectively. It is therein shown that after a fast transient period, the initially smooth data form one left and one right going shock waves, which are followed by rarefactions spreading towards the opposite direction of the shock wave movement. In fact, $h$ is symmetric along the spatial domain, while $q$ is antisymmetric; see Figure A.2c. From the 4 numerical solutions of the SW equations, we randomly split the 15-15-70% of the time-series data to form the training/validation/testing data sets of the GRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=0.5$ and $\sigma=0.4$ are denoted with black circles in Figure A.2a,b.
@@
Figure A.2: Numerical solution of the SW (depth, momentum) and PW (density, momentum) equations obtained with the high-resolution Godunov-scheme in section 3.1 for (a,b,c) $x\in[-5,5]$, $t\in[0,3]$, with the initial data $h(0,x)=0.5~{}exp(-x^{2}/(2\cdot 0.4^{2}))$ and $q(0,x)=0$ shown in panel (c), and (d,e,f) $x\in[0,800]$, $t\in[0,600]$, with the initial data $\rho(0,x)=0.1~{}(1+0.3sin(2\pi x/L))$ and $q(0,x)=0.1$ shown in panel (f). Periodic boundary conditions are used for both SW and PW equations. The black circles on (a,b,c,d) denote the time steps $n_{t}$, collected from the depicted numerical solutions, to form the training set of GRINNs.
@@ A.4 The PW equations of traffic flow
For the derivation of $\hat{\mathbf{A}}_{i-1/2}$, we followed the general approach introduced by [66], as presented in detail in section C.4. The resulting numerical solution of the PW equations, using the Van-Leer flux-limiter function, for the sinusoidal initial data with $\mu=0.3$ is depicted in Figure A.2d,e,f for $\rho$ and $q$, respectively. It is therein shown that the initial pulse gradually forms a left going shock wave, shown by the abrupt difference of low (high) values of $\rho$ ($q$) to high (low) ones. Trailing to the right of the travelling shock wave, fluctuations develop which after $t>500$ lead to the development of another left going shock, with the opposite behavior; abrupt differences from high (low) values of $\rho$ ($q$) to low (high) ones. From the 4 numerical solutions of the PW equations, we randomly split the 7.5-7.5-85% of the time-series data to form the training/validation/testing data sets of the GRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=0.3$ are denoted with black circles in Figure A.2d,e.
@@ Appendix D Supplemental Figures for PW equation when learning $\mathcal{N}(\rho,q)$
(d) $P(\rho)$ vs GRINNs $\mathcal{N}(\rho,q)$
Figure D.1: Numerical accuracy of GRINNs, assuming a pressure closure of the form $\mathcal{N}(\rho,q)$, for the PW equations. (a,b) Absolute errors of the numerical solution provided by the GRINNs learned equations in eq. 26 ($\hat{\rho}$ and $\hat{q}$) vs the one provided by the PW equations in eq. A.6 ($\rho$ and $q$). (c) Solutions compared in panel (a) at $t=500$. (d) Analytically known pressure closure $P(\rho)$ vs the $\mathcal{N}(\rho,q)$ functional learned with GRINNs.

### v2

@@ 5 Conclusions
GoRINNs offer interpretable shallow feedforward NNs for tackling the above problems. Unlike other hybrid SciML methods that approximate the numerical fluxes, GoRINNs learn the solution of the Godunov scheme by directly approximating the physical flux functions per se. This approach enhances interpretability, thus respecting explicitly the underlying conservation laws as approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. Moreover, GoRINNs provide a versatile framework for identifying flux functionals directly from data. This approach accommodates systems of conservation laws and allows for the inclusion of source terms—whether known or unknown—in the associated hyperbolic PDEs. As we show, GoRINNs are able to approximate accurately the closures of four benchmark problems, which exhibit traveling shocks that emerge at finite times. Based on the solution of the inverse problem, which actually approximates the physical flux function required for the solution operator of the Godunov-type scheme, we solved the forward problem for long time periods, showing that GoRINNs result to a high numerical approximation accuracy, preserving numerical stability and the conservation laws.
The training and performance of GoRINNs, like other SciML methods, benefit from partial knowledge of the physics underlying the missing terms (see section 4.4, where GoRINNs converge with the “correct” predictor selection.) If no physical insight is available, then techniques such as symbolic regression [75, 6], manifold learning (e.g., parsimonious diffusion maps) [15, 22, 20, 25, 56], and SINDy [8, 10] can be exploited to learn a parsimonious set of variables that can parametrize the unknown closures. Furthermore, GoRINNs can be extended to analyze real-world data derived from microscopic simulations/observations, facilitating the discovery of macroscopic hyperbolic PDEs for the emergent dynamics. This application will further require parameter inference, a task that can be addressed by including physical parameters as additional GoRINNs predictors. Finally, using GoRINNs to discover macroscopic closures connects directly to control theory, which can be exploited for enabling distributed control in large-scale microscopic models (see e.g., [47, 48]), given GoRINNs-derived information at the macroscopic level.
@@ Appendix A Benchmark problems and data acquisition
Here, we present the four benchmark problems upon which the performance of GoRINNs is evaluated on, namely the inviscid Burgers’ equation, the LWR equation of traffic flow, the SW equations of fluid dynamics, and the PW equations of traffic flow. We additionally provide the details for the solution of the forward problem with the high-resolution Godunov-type FV scheme in section 3.1, which are required for the collection of training/validation/testing data sets.
@@ A.1 Burger’s equation
In spite of the smooth initial conditions, contact discontinuities are formed within the selected time interval, in particular shock waves and rarefactions; the latter are not transonic due to the positive initial conditions. Thus, for the numerical solution of the forward problem via the Godunov-scheme in section 3.1, it is sufficient to employ the Roe solver, which leads to the scalar, in this case, Roe matrix $\hat{a}_{i-1/2}=\bar{u}$, where $\bar{u}=(u_{l}+u_{r})/2$; the derivation is provided in section C.1. As already discussed in section 3.3, the Van-Leer flux-limiter function is employed. The resulting numerical solution of the Burgers’ equation for the Gaussian initial data with $\mu=2$ and $\sigma=0.2$ is depicted in Figure A.1a,b, where it is clearly shown that the initial bump quickly forms a right going shock wave, followed by a rarefaction spread out towards the left; for comparison to the analytic solution, see Chapter 4 of [35]. From the 4 numerical solutions of the Burgers’ equation, we randomly split the 15-15-70% of the time-series data to form the training/validation/testing data sets of the GoRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=2$ and $\sigma=0.2$ are denoted with black circles in Figure A.1a.
@@
Figure A.1: Numerical solution of the Burgers’ and LWR equations obtained with the high-resolution Godunov-scheme in section 3.1 for (a,b) $x\in[-1,1]$, $t\in[0,3]$, with the initial data $u(0,x)=2~{}exp(-x^{2}/(2\cdot 0.2^{2}))$ shown in panel (b), and (c,d) $x\in[-20,20]$, $t\in[0,60]$ with the $\rho(0,x)=exp(-(x+10)^{2}/(2\cdot 2^{2}))$ shown in panel (d). Periodic boundary conditions are used for the Burgers’ equation and outflow ones for the LWR equation. The black circles on (a,c) denote the time steps $n_{t}$, collected from the depicted numerical solutions, to form the training set of GoRINNs.
@@ A.2 LWR equation of traffic flow
Similarly to the Burgers’ equation, in spite of the smooth initial conditions, contact discontinuities are formed within the selected time interval. For the numerical solution via the Godunov-scheme in section 3.1, it is sufficient to employ the Roe solver. The scalar, for the LWR equation, Roe matrix takes the form $\hat{a}_{i-1/2}=v_{max}(1-2\bar{\rho})$, where $\bar{\rho}=(\rho_{l}+\rho_{r})/2$; see the derivation in section C.2. The resulting numerical solution of the LWR equation, using the Van-Leer flux-limiter function, for the Gaussian initial conditions with $\sigma=2$ is depicted in Figure A.1c,d. In contrast to the Burgers’ equation, the non-linear flux term of the LWR equation $f(\rho)=v_{max}\rho(1-\rho)$ is not convex, but concave. Because of that, as shown in Figure A.1c, the initial smooth profiles quickly form a right going shock wave, which is now following the right-spreading rarefaction (instead of preceding it, as shown in Figure A.1a for the Burgers’ equation). This behavior is well aligned with the relaxation of traffic jams reported for agent-based car-following models; see Chapter 11 of [41] for a direct comparison. From the 4 numerical solutions of the LWR equation, we randomly split the 15-15-70% of the time-series data to form the training/validation/testing data sets of the GoRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=1$ and $\sigma=2$ are denoted with black circles in Figure A.1c.
@@ A.3 The SW equations
the latter being the so-called Roe average. For the derivation of $\hat{\mathbf{A}}_{i-1/2}$, which is presented in detail in section C.3, we followed the general approach introduced by [66], resulting in the same Roe matrix as the ones derived in [41, 35]. The resulting numerical solution of the SW equations, using the Van-Leer flux-limiter function, for the Gaussian initial data with $\sigma=0.4$ is depicted in Figure A.2a,b,c for $h$ and $q$, respectively. It is therein shown that after a fast transient period, the initially smooth data form one left and one right going shock waves, which are followed by rarefactions spreading towards the opposite direction of the shock wave movement. In fact, $h$ is symmetric along the spatial domain, while $q$ is antisymmetric; see Figure A.2c. From the 4 numerical solutions of the SW equations, we randomly split the 15-15-70% of the time-series data to form the training/validation/testing data sets of the GoRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=0.5$ and $\sigma=0.4$ are denoted with black circles in Figure A.2a,b.
@@
Figure A.2: Numerical solution of the SW (depth, momentum) and PW (density, momentum) equations obtained with the high-resolution Godunov-scheme in section 3.1 for (a,b,c) $x\in[-5,5]$, $t\in[0,3]$, with the initial data $h(0,x)=0.5~{}exp(-x^{2}/(2\cdot 0.4^{2}))$ and $q(0,x)=0$ shown in panel (c), and (d,e,f) $x\in[0,800]$, $t\in[0,600]$, with the initial data $\rho(0,x)=0.1~{}(1+0.3sin(2\pi x/L))$ and $q(0,x)=0.1$ shown in panel (f). Periodic boundary conditions are used for both SW and PW equations. The black circles on (a,b,c,d) denote the time steps $n_{t}$, collected from the depicted numerical solutions, to form the training set of GoRINNs.
@@ A.4 The PW equations of traffic flow
For the derivation of $\hat{\mathbf{A}}_{i-1/2}$, we followed the general approach introduced by [66], as presented in detail in section C.4. The resulting numerical solution of the PW equations, using the Van-Leer flux-limiter function, for the sinusoidal initial data with $\mu=0.3$ is depicted in Figure A.2d,e,f for $\rho$ and $q$, respectively. It is therein shown that the initial pulse gradually forms a left going shock wave, shown by the abrupt difference of low (high) values of $\rho$ ($q$) to high (low) ones. Trailing to the right of the travelling shock wave, fluctuations develop which after $t>500$ lead to the development of another left going shock, with the opposite behavior; abrupt differences from high (low) values of $\rho$ ($q$) to low (high) ones. From the 4 numerical solutions of the PW equations, we randomly split the 7.5-7.5-85% of the time-series data to form the training/validation/testing data sets of the GoRINNs, as discussed in section 3.3; the training samples collected from the solution with initial data with $\mu=0.3$ are denoted with black circles in Figure A.2d,e.
@@ Appendix D Supplemental Figures for PW equation when learning $\mathcal{N}(\rho,q)$
(d) $P(\rho)$ vs GoRINNs $\mathcal{N}(\rho,q)$
Figure D.1: Numerical accuracy of GoRINNs, assuming a pressure closure of the form $\mathcal{N}(\rho,q)$, for the PW equations. (a,b) Absolute errors of the numerical solution provided by the GoRINNs learned equations in eq. 26 ($\hat{\rho}$ and $\hat{q}$) vs the one provided by the PW equations in eq. A.6 ($\rho$ and $q$). (c) Solutions compared in panel (a) at $t=500$. (d) Analytically known pressure closure $P(\rho)$ vs the $\mathcal{N}(\rho,q)$ functional learned with GoRINNs.